### v1

@@ 4.2 Iterative Diffusive: Recursive Algorithm for Heterogeneous Systems
As in the original scheme, the algorithm proceeds in discrete steps, where all existing processes collaborate to spawn new groups on additional nodes. The assignment of a unique group identifier (group_idgroup\_id), ranging from 0 to the total number of groups to be created, is also preserved.
@@
Three vectors are needed to describe the system configuration. The number of entries in these vectors is equal to the number of nodes in the system (N).
  • $M$: Contains the number of available cores assigned to a job on each node.
  • AA: Stores the number of processes running on each node before starting this algorithm.
@@
  • $S$: Has the number of processes to be created in each node when executing this algorithm.
@@
Based on these vectors, a stepwise sequence can be constructed to determine the total number of processes and nodes involved at each spawning step. Only nodes with positive entries in the vector $S$ will allocate new processes; thus, each non-zero entry indicates the creation of a new group on the corresponding node. This vector is computed from the previous ones as follows:
$S_{i}=M_{i}-A_{i},\quad\text{for }i=0,1,\dots,N-1$
@@
Equation 4 defines the total number of processes $t_{s}$ present in step $s$, as a function of the cumulative total of previous steps ($t_{s-1}$) and the number of processes generated in the current step ($g_{s}$). Step $s=0$ corresponds to the initial state, which includes only the source processes, as specified by the vector $A$.
$t_{s}=\begin{cases}\sum\limits_{j=0}^{N-1}A_{j},&\text{if }s=0\\ t_{s-1}+g_{s},&\text{if }s>0\end{cases}$ (4)
Equation 5 defines the number of processes $g_{s}$ generated in step $s$. At each stage, the elements of vector $S$ from index $\nu_{s}$ to index $\lambda_{s}$ are accumulated to calculate the number of new processes that need to be generated.
$g_{s}$ $=\sum\limits_{j=\nu_{s}}^{\lambda_{s}}S_{j}$ $\quad\text{where}\quad\lambda_{s}$ $=\min\left(N-1,\,\left(\sum\limits_{i=0}^{s-1}t_{i}\right)-1\right),$ (5)
$\nu_{s}$ $=\sum\limits_{i=0}^{s-2}t_{i}$
Finally, Equation 7 represents the cumulative number of occupied nodes ($T_{s}$) up to step $s$, whereas Equation 6 specifies the number of new nodes ($G_{s}$) occupied by newly created groups during that step. In all equations, null elements in vector $S$ are disregarded.
$G_{s}=\min(N-T_{s-1},\,t_{s-1})$ (6)
$T_{s}=\begin{cases}I,&\text{if }s=0\\ T_{s-1}+G_{s},&\text{if }s>0\end{cases}$ (7)
@@ 4.3 Synchronization Between Process Groups
Once all processes have been created, it is necessary to establish connections between them to enable direct communication. With the adopted creation strategy, each newly spawned group initially knows only its parent process (the one that created it) and, if applicable, its own child processes. As shown in Figure 1, processes can only communicate directly with those to which they are connected by an edge in the spawn graph.
@@
First, the process groups must be prepared for connection. The processes designated as roots in the MPI_Comm_accept calls must first open a port using MPI_Open_port and publish the service name via MPI_Publish_name. Since MPI_Comm_connect must be called after the corresponding port has been opened by the other group, a global synchronization phase between all groups is required. Otherwise, execution errors may occur, as observed with implementations such as MPICH [^1] This behaviour has only been tested with MPICH..
The synchronization between groups consists of three steps and is carried out within each group using a dedicated subcommunicator. The objective is to ensure that all processes are aware of the port status before attempting to establish a connection.
  1. [1]
  2. 1.
@@
    Subcommunicator creation. Each group creates a subcommunicator using MPI_Comm_split, including the root process of the group and all processes that have created child groups. This subcommunicator will be used to coordinate intra-group synchronization.
  3. 2.
@@
    Upside Synchronization. Each process that has spawned child groups waits to receive a token from each of them using MPI_Irecv followed by MPI_Waitall. Then, all processes in the subcommunicator execute a call to MPI_Barrier, ensuring that all tokens have been received. Finally, the root process (if not part of the source group) notifies its parent group using MPI_Send.
  4. 3.
@@
    Downside Synchronization The root process of each group (except the source group) waits to receive a token from its creator via MPI_Recv. Then, processes in the subcommunicator perform a MPI_Barrier call (excluding the source group). Finally, each process that has spawned child groups sends them a token using MPI_Isend followed by MPI_Waitall. These barriers are essential to ensure that no group proceeds before all of its descendants have completed the Upside step, and to prevent any group from notifying its children before being notified by its parent group.
Listing 1 shows the implementation of the synchronization mechanism. The processes use the world_c communicator, which corresponds to MCW for spawned processes, or to a shared communicator among source processes in the case of creators. The parent_c communicator is an intercommunicator between a spawned group and its parent. The variable qty_s indicates the number of child groups spawned by a process, and the corresponding intercommunicators are stored in the spawn_c vector.
@@
2 int qty_s, MPI_Comm *spawn_c) {
3 int myId, i, root = 0;
4 int color = qty_s ? 1 : MPI_UNDEFINED;
5 MPI_Request *reqs = malloc(qty_s * sizeof *reqs);
@@ 6 MPI_Comm synch_ranks;
9 MPI_Comm_rank(world_c, &myId);
10 MPI_Comm_split(world_c, color, myId, &synch_ranks);
11
@@ 10 MPI_Comm_split(world_c, color, myId, &synch_ranks);
13 for(i=0; i<qty_s; i++) {
14 MPI_Irecv(…, source=root, spawn_c[i], &reqs[i]);
@@
16 if(qty_s){MPI_Waitall(reqs);MPI_Barrier(synch_ranks);}
17
18 if(parent_c != MPI_COMM_NULL && myId == root) {
19 // Send to parent this group is ready
20 MPI_Send(…, dest=root, parent_c);
@@
22 MPI_Recv(…, source=root, parent_c);
@@
24
25 // Notify this group that all groups are ready
26 if(parent_c != MPI_COMM_NULL && qty_s)
@@
35}
Listing 1: Syncrhonization step before binary connection.
Figure 2 illustrates the upside synchronization and downside synchronization steps, shown on the left and right sides, respectively. In this example, six process groups are created in two spawning steps. The arrows indicate which processes send messages to signal their readiness in the upside phase, and to confirm that all groups are ready in the downside phase. A group represented in black with stripes indicates that it must wait to receive all messages before proceeding. At that point, a MPI_Barrier is used to synchronize the group, after which the execution can continue.
Figure 2: Synchronization of 6 spawned groups and the initial node. The arrows indicate the direction of messages, and stripes the need of that group of using a Barrier.
@@ 4.4 Binary connection
Once synchronization between all groups is complete, the connection operation can begin. The adopted strategy consists of making pairwise binary connections in several successive steps. In each step, groups with an identifier less than half the number of active groups perform an MPI_Comm_accept operation, while those with a higher identifier execute an MPI_Comm_connect. In the case of an odd number of groups, the middle group is temporarily excluded from that step and proceeds directly to the next one.
Each group connects to the group whose identifier is its mirror image within the current set of active groups, that is, the group symmetrically positioned relative to the center. Once the connection is completed, both groups are merged into a new communicator and adopt a common group identifier, corresponding to the group that executed the MPI_Comm_accept call.
@@
Figure 3: Connection of 7 spawned groups in 3 steps. In each step the amount of groups is halved and connections operations may be unordered.
Listing 2 presents a simplified version of the code used to implement this binary connection strategy between groups. In the code, the groups are distinguished based on whether they should perform a MPI_Comm_accept operation (lines L11–L15) or a MPI_Comm_connect (lines L17–L23). In the latter case, the group obtains the corresponding port name (line L19) using the group_id of the target group and the MPI_Lookup_name function.
@@
3 int middle, new_groups, new_gid;
@@ 5 MPI_Comm merge_comm, aux_comm, intercomm;
10 new_groups = ceil(groups / 2.0);
@@
12 MPI_Comm_accept(my_port, MPI_INFO_NULL,
13 MAM_ROOT, merge_comm, &intercomm);
@@ 14 MPI_Intercomm_merge(intercomm, 0, &aux_comm);
20 MPI_Comm_connect(port, MPI_INFO_NULL,
21 MAM_ROOT, merge_comm, &intercomm);
@@ 22 MPI_Intercomm_merge(intercomm, 1, &aux_comm);
27}
Listing 2: Binnary connection step for the parallel spawn.
4.5 Rank Reordering and Connection with sources
@@
Once all the generated groups have been merged into a single one, the process ranks may not be ordered across nodes. This is because binary connections between groups do not enforce a specific order and are susceptible to race conditions.
@@
To ensure a consistent global order, an MPI_Comm_split call is made, where each process specifies its new rank as follows:
$\text{world\_rank}+\sum\limits_{j=0}^{N-1}A_{j}+\sum\limits_{j=0}^{\text{group\_id-1}}S_{j}$
At the same time, all intermediate communicators created throughout the strategy are disconnected using MPI_Comm_disconnect. This frees up resources and ensures that any node can be removed or replaced in future reconfigurations without leaving residual dependencies.
Finally, the target group is connected to the source group, establishing the final inter-communicator that links them. This operation concludes the process creation and connection phases.

### v2

@@ 4.2 Iterative Diffusive Strategy
As in the previous scheme, the algorithm proceeds in discrete steps, where all existing processes collaborate to spawn new groups on additional nodes. The assignment of a unique group identifier (group_idgroup\_id), ranging from 0 to the total number of groups to be spawned, is also preserved.
@@
Three vectors are needed to describe the allocation configuration, and are maintained through the execution. The number of entries in these vectors is equal to the number of nodes in the allocation ($N$).
  • AA: Contains the total number of cores assigned to the job on each node.
  • $R$: Stores the number of processes of the job running on each node.
@@
  • $S$: Keeps the number of processes of the job to be spawned in each node when executing this algorithm. This vector is computed from the previous ones as follows:
$S_{i}=A_{i}-R_{i},\quad\text{for }i=0,1,\dots,N-1.$
@@
Based on these vectors, a stepwise sequence can be constructed to determine the total number of processes and nodes involved at each spawning step. Note that only nodes with positive entries in the vector $S$ will allocate new processes; thus, each entry greater than zero indicates the creation of a new group on the corresponding node. The number of existing processes in each step $s$ can then be calculated as follows:
$t_{s}=\begin{cases}\sum\limits_{j=0}^{N-1}R_{j},&\text{if }s=0\\ t_{s-1}+g_{s},&\text{if }s>0\end{cases}\quad,$ (4)
@@
where $t_{s}$ is the total processes existing at the end of step $s$, calculated as a function of the cumulative total of previous steps ($t_{s-1}$) and the number of processes generated in the current step ($g_{s}$). The initial state ($s=0$) corresponds to the number of source processes specified in vector $R$, and the number of generated processes ($g_{s}$) in a given step can be calculated as the sum of a sequence of elements in vector $S$:
$g_{s}=\sum\limits_{i=\lambda_{s-1}}^{\min\left(N,\,\lambda_{s}\right)-1}S_{i},$ (5)
where the limits $\lambda_{s}$ defines the range of indices of vector $S$ used at each spawning step, as follows:
$\lambda_{s}=\begin{cases}0,&\text{if }s=0\\ \lambda_{s-1}+t_{s-1},&\text{if }s>0\end{cases}\quad.$ (6)
These limits ensure that each step consumes a contiguous segment of $S$, without overlap or omission, while maintaining a sequential progression through the vector.
Specifically, $\lambda_{s-1}$ marks the starting index for step $s$, which corresponds to the total number of processes handled in all previous steps. Conversely, $\lambda_{s}-1$ marks the last index to be included in the current step, extending over the $t_{s-1}$ processes that participate at this stage. As a result, each step $s$ uses exactly $t_{s-1}$ entries of $S$, starting right after the portion used in step s1s-1. The minimum in the upper index for Equation 5 is used to ensure no more than $N$ spawns are performed. In all equations, null elements in vector $S$ are disregarded.
The cumulative number of occupied nodes ($T_{s}$) up to step $s$ is calculated from these previous equations as follows:
$T_{s}=\begin{cases}I,&\text{if }s=0\\ T_{s-1}+G_{s},&\text{if }s>0\end{cases}\quad,$ (7)
where for each step, the sum of the nodes from the previous step ($T_{s-1}$) and the nodes added during the current step ($G_{s}$) are calculated. In the first step ($s=0$) the initial number of nodes is used ($I$). For the next steps the number of nodes mainly depends on $G_{s}$, which represents the number of new nodes to add. A new node is added if the corresponding positions in the vectors $R$ and $S$ do not contain any processes before spawning ($R_{i}=0$) and if there are processes to spawn ($S_{i}>0$). Therefore, one node is added for each position considered in the range defined by Equation 6, only if both conditions hold. This can be represented as follows:
$G_{s}=\sum\limits_{i=\lambda_{s-1}}^{\min\left(N,\,\lambda_{s}\right)-1}\{R_{i}=0\wedge S_{i}>0\},$ (8)
where the limits $\lambda_{s}$ are calculated by using Equation 6. When the condition of the summation is met, it adds 1, or 0 otherwise.
Table 2 shows how the different Equations behave depending on the initial parameters $A$, $R$, $S$, NN and $I$. In this example, $s=0$ is the initial state before executing this strategy.
Initial Vectors
A [4, 2, 8, 12, 3, 3, 4, 4, 6, 3]
R [2, 0, 0,  0,  0, 0, 0, 0, 0, 0]
S [2, 2, 8, 12, 3, 3, 4, 4, 6, 3]
Computed Values per Step
$s$ $t_{s}$ $g_{s}$ $\lambda_{s}$ $T_{s}$ $G_{s}$
0 2 - 0 1 -
1 6 4 2 2 1
2 40 34 7 8 6
3 49 9 47 10 2
Table 2: Iterative diffusive procedure example for a new allocation of resources from 11 node (II) to $10$ nodes ($N$).
Similar to the previous approach, this one ensures that no shared MCW exists across the expanded nodes, with the added advantage of accounting for a heterogeneous number of cores on each expanded node. Moreover, it also enables fast shrinkage through TS method.
@@ 4.3 Synchronization Between Process Groups
The purpose of synchronizing the groups is to ensure that all processes know that every port is active before any connection attempt is made. Once all processes have been spawned, establishing these connections is essential, as they enable direct communication among them. Due to the chosen process creation strategies, each newly spawned group only knows its parent process (its creator) and, if applicable, its own child processes. Consequently, as shown in Figure 1, processes can only communicate directly with those linked to them by an edge in the spawn graph.
@@
First, the process groups must be prepared for connection. Thus, the processes designated as roots in the MPI_Comm_accept calls must first open a port using MPI_Open_port and publish the service name via MPI_Publish_name. Since MPI_Comm_connect must be called after the corresponding port has been opened by the other group, a global synchronization phase between all groups is required. Otherwise, execution errors may occur, as observed with implementations such as MPICH [^1] This behavior has only been tested with MPICH.. For the rest of the paper, the root process of each group is always the process with rank 0 in its communicator, which is required by the functions discussed here.
Each group carries out the synchronization using a dedicated subcommunicator consisting of three stages.
  1. 1
@@
    Subcommunicator creation. Each group creates a subcommunicator using MPI_Comm_split, including the root process of the group and all processes of the group that have spawned child groups. This subcommunicator will be used to coordinate intra-group synchronization.
  2. 2
@@
    Upside Synchronization. Each process that has spawned child groups waits to receive a token from each of them using MPI_Irecv and MPI_Waitall. Then, all processes in the subcommunicator execute a call to MPI_Barrier, ensuring that all tokens have been received. Finally, the root process (if not part of the source group) notifies its parent group using MPI_Send.
  3. 3
@@
    Downside Synchronization. The root process of each group (except the source group) waits to receive a token from its creator via MPI_Recv. Then, processes in the subcommunicator perform a MPI_Barrier call (excluding the source group). Finally, each process that has spawned child groups sends them a token using MPI_Isend, followed by MPI_Waitall.
The barriers included in the stages are essential to ensure that no group proceeds before all its descendants have completed the Upside step, and to prevent any group from notifying its children before being notified by its parent group. However, the Waitall is required only to ensure the communication is completed.
The left and right sides of Figure 2 illustrate, respectively, the upside synchronization and downside synchronization steps. In this example, six process groups are spawned in two spawning steps. The arrows indicate which processes send messages to signal their readiness during the upside phase, and to confirm that all groups are ready in the downside phase. A group represented in black with stripes indicates that it must wait to receive all messages before proceeding. At that point, a MPI_Barrier is used to synchronize the group, after which the execution can continue.
Listing 1 shows the implementation of the synchronization mechanism, where the first stage is shown in lines L13-L17, the second from lines L19-L28, and the third from lines L30-L41. The processes use the world_c communicator, which corresponds to MCW for spawned processes or to the global communicator among source processes. Then, the parent_c communicator is an inter-communicator between a spawned group and its parent. Additionally, the variable qty_c indicates the number of child groups spawned by the calling process, and the corresponding inter-communicators are stored in the spawn_c vector. A particular MPI_Barrier function is needed at line L24, because other ranks of the same group could also have children, and therefore the group has to ensure all messages have been received before continuing. This is represented for the Init Node group and the Group_Id 0 group in the upside synchronization in Figure 2.
@@
2 int qty_c, MPI_Comm *spawn_c) {
3 // world_c: Built comm for sources, MCW for targets
4 // parent_c: Intercomm to contact parent rank if any
5 // qty_c: Total children groups for calling rank
6 // spawn_c: Children comms array for calling rank
7 int rank, i, root = 0;
8 // The color separates ranks into two subgroups
9 int color = qty_c ? 1 : MPI_UNDEFINED;
10 MPI_Request *reqs = malloc(qty_c * sizeof(MPI_Request));
@@ 11 MPI_Comm synch_ranks;
15 // synch_ranks is only used by ranks with children
16 MPI_Comm_rank(world_c, &rank);
17 MPI_Comm_split(world_c, color, rank, &synch_ranks);
18
19// Stage 2: Upside Synchronization
@@ 17 MPI_Comm_split(world_c, color, rank, &synch_ranks);
21 for(i=0; i<qty_c; i++) {
22 MPI_Irecv(…, root, spawn_c[i], &reqs[i]);
@@
24 if(qty_c){ MPI_Waitall(reqs); MPI_Barrier(synch_ranks); }
25 if(parent_c != MPI_COMM_NULL && rank == root) {
26 // Send to parent that this group is ready
27 MPI_Send(…, root, parent_c);
28 }
29
30// Stage 3: Downside Synchronization
31 if(parent_c != MPI_COMM_NULL && rank == root) {
@@ 27 MPI_Send(…, root, parent_c);
33 MPI_Recv(…, root, parent_c);
@@
35 // Notify all ranks in the group to continue
36 if(parent_c != MPI_COMM_NULL && qty_c)
@@
45}
Listing 1: ]Synchronization step before binary connection.
Figure 2: Synchronization of $6$ spawned groups and the initial node. The arrows indicate the direction of messages, and the stripes the necessity of using Barriers.
@@ 4.4 Binary connection
Once synchronization between all groups is completed, the connection operation can begin. The adopted strategy consists of establishing binary connections between pairs of groups in several successive steps. In each one, groups with an identifier less than half the number of active groups perform an MPI_Comm_accept, while those with a higher identifier execute an MPI_Comm_connect. In the case of an odd number of groups, the middle group is temporarily excluded from that step and proceeds directly to the next one.
Once the connection is completed, both groups are merged into a new communicator and adopt a common group identifier, corresponding to the group that executed the MPI_Comm_accept call.
@@
Figure 3: Connection of $7$ spawned groups in 33 steps. In each step the number of groups is halved and connections operations may be unordered.
Listing 2 presents a simplified version of the code used to implement this binary connection strategy between groups. In the code, the groups are distinguished based on whether they should perform an MPI_Comm_accept operation (lines L15–L19) or an MPI_Comm_connect (lines L20–L26). In the latter case, the group obtains the corresponding port name by calling the get_remote_port function (line L22), which uses the target group’s group_id to return the port name. To do so, it first writes an established service name based on the group_id, and then calls MPI_Lookup_name, which retrieves the port name associated with the service.
@@
3 // groups: Total spawned groups.
4 // group_id: Group identifier for calling rank
5 // my_port: Local port name of calling rank if any
6 // new_comm: Output communicator of combined groups
7 int middle, new_groups, new_gid, root = 0;
@@ 9 MPI_Comm merge_comm, aux_comm, intercomm;
14 new_groups = groups - middle;
@@
16 MPI_Comm_accept(my_port, MPI_INFO_NULL, root,
17 merge_comm, &intercomm);
@@ 18 MPI_Intercomm_merge(intercomm, 0, &aux_comm);
23 MPI_Comm_connect(port, MPI_INFO_NULL, root,
24 merge_comm, &intercomm);
@@ 25 MPI_Intercomm_merge(intercomm, 1, &aux_comm);
31}
Listing 2: ]Binary connection step for the parallel spawn.
4.5 Rank Reordering
@@
Once all the generated groups have been merged into a single one, the MPI ranks may not be ordered across nodes. This is because binary connections between groups do not enforce a specific order and, therefore, are susceptible to race conditions.
@@
To ensure a consistent global order, an MPI_Comm_split call is made in which each process specifies its new rank as follows:
$\text{world\_rank}+\sum\limits_{j=0}^{N-1}R_{j}+\sum\limits_{j=0}^{\text{group\_id-1}}S_{j},$ (9)
where, world_rank is the rank that the calling process had in its MCW, while both summations take into account all the previous ranks. The first one counts the number of ranks that existed before the resize, but is not computed since the value is obtained when the group is spawned and can be used directly. Meanwhile, the second summation counts the number of ranks in all groups whose group_id is smaller than that of the calling rank.
4.6 Overview
To conclude this subsection, a global overview of the tasks required to perform the parallel spawn is provided, both for source and newly spawned processes.
The steps for sources must be carried out in the following order:
  1. 1.
    Open a port and service name for the root rank.
  2. 2.
    Use one of the strategies described in Subsections 4.1 and 4.2 to perform the parallel spawn.
  3. 3.
    Synchronize ranks to ensure correctness during the binary connection. (Subsection 4.3)
  4. 4.
    Connect the merged-groups communicator with the initial group communicator.
Listing 3 shows these tasks for the source processes. First, they compute the required initial data (line L14), and then the root process opens a port (line L17). At this point, the parallel spawn begins, selecting the appropriate strategy depending on whether the new node allocation is homogeneous or heterogeneous (lines L20–L24). Note that, each algorithm determines whether a given rank must spawn additional groups. Next, the ranks synchronize to ensure correctness (line L27). Afterwards, unnecessary communicators are disconnected (lines L30-L31), and finally the initial group waits for the merged group of spawned processes to connect to the port opened by the root process (lines L32-L34).
In the case of newly spawned processes, the steps to carry out are the following:
  1. 1.
    Open a port for root ranks whose group identifier satisfies that group_id<total_spawned_groups/2{group\_id}<{total\_spawned\_groups}/2. (Prior preparation to Subsection 4.4)
  2. 2.
    Use one of the strategies described in Subsections 4.1 and 4.2 to perform the parallel spawn.
  3. 3.
    Synchronize ranks to ensure correctness during the binary connection. (Subsection 4.3)
  4. 4.
    Carry out the binary connection to merge the spawned groups into a single one. (Subsection 4.4)
  5. 5.
    Reorder the ranks to match the expected ordering. (Subsection 4.5)
  6. 6.
    Connect the merged-groups communicator with the initial group communicator.
Listing 4 shows the tasks carried out for the generated groups. At startup, they first obtain the data they will use, which they receive from their parent group upon creation (lines L17-L18). Then, in each group, the root rank of those groups whose group_id is lower than half of the generated groups (groups–I) proceeds to open a port (lines L21–L22). At this point, the parallel spawn begins, selecting the appropriate strategy depending on whether the new node allocation is homogeneous or heterogeneous (lines L25–L29). Once all groups have been created, they synchronize to ensure that all ports have been opened (line L32), and then disconnect any communicators that are no longer needed (lines L33–L36). After this synchronization, all spawned processes are merged into a single communicator (lines L39-L40) and then reordered according to the expected rank ordering (line L41). Finally, the root process of the merged group obtains the port name of the initial group (lines L44-L45) and connects to it (lines L46-L47).
1void parallel_strat_parents(MPI_Comm group_comm, int rank,
2 MPI_Comm *out_comm) {
3 // group_comm: Built comm for sources
4 // rank: Identifier of calling process
5 // out_comm: New intercomm for sources and spawned
6 int qty_c;
7 int groups, root;
8 int N, *A, *R, *S;
9 char *port_name;
10 MPI_Comm *spawn_c = NULL;
11
12 qty_c = 0; root = 0;
13 // Calculate following data
14 get_initial_data(&N,&A,&R,&S, &groups);
15
16 //Open port if rank 0 (root)
17 open_port(&port_name, !rank, groups);
18
19 // Decide spawn depending on core counts
20 if(check_homogenous_dist(S)) {
21 hypercube(…, &qty_c, &spawn_c);
22 } else {
23 diffusive_iterative(…, &qty_c, &spawn_c);
24 }
25
26 // Synchronize all ranks for the binary connection
27 common_synch(group_comm, MPI_COMM_NULL, qty_c, spawn_c);
28
29 // Create intercomm between sources and children
30 for(int i=0; i<qty_c; i++)
31 {MPI_Comm_disconnect(&spawn_c[i]);}
32 if(spawn_c != NULL) free(spawn_c);
33 MPI_Comm_accept(port_name, MPI_INFO_NULL, root,
34 group_comm, out_comm);
35}
Listing 3: ]Overall tasks for source ranks.
1void parallel_strat_children(int rank, MPI_Comm *out_comm,
2 MPI_Comm *newintracomm) {
3 // rank: Identifier of calling process
4 // out_comm: New intercomm for sources and spawned
5 // newintracomm: Output intracomm for spawned ranks
6 int i, exp_id, group_id, opening, qty_c;
7 int groups, root;
8 int N, I, *A, *R, *S;
9 int initial_group_id = -1;
10 char *port_name, *remote_port_name;
11 MPI_Comm *spawn_c = NULL;
12 MPI_Comm group_comm = MPI_COMM_WORLD;
13 MPI_Comm_get_parent(out_comm);
14
15 qty_c = 0; root = 0;
16 // Data is obtained when rank is spawned
17 get_initial_data_children(&N,&A,&R,&S, &groups, &I,
18 &group_id);
19
20 // Open port if required for binary connection
21 opening = (rank == root && group_id < (groups-I)/2) ?1:0;
22 open_port(&port_name, opening, group_id);
23
24 // Decide spawn depending on core counts
25 if(check_homogenous_dist(S)) {
26 hypercube_spawn(…, &qty_c, &spawn_c);
27 } else {
28 diffusive_iterative_spawn(…, &qty_c, &spawn_c);
29 }
30
31 // Synchronize all ranks for the binary connection
32 common_synch(group_comm, *out_comm, qty_c, spawn_c);
33 for(i=0; i<qty_c; i++)
34 { MPI_Comm_disconnect(&spawn_c[i]); }
35 if(spawn_c != NULL) free(spawn_c);
36 MPI_Comm_disconnect(out_comm);
37
38 // Connect groups and ensure expected rank order
39 binary_connection(groups - I, group_id, port_name,
40 newintracomm);
41 rank_reorder(newintracomm, N, R, S, rank);
42
43 // Create intercomm between sources and children
44 if(rank == root)
45 get_remote_port(initial_group_id, &remote_port_name);
46 MPI_Comm_connect(remote_port_name, MPI_INFO_NULL, root,
47 *newintracomm, out_comm);
48}
Listing 4: ]Overall tasks for target ranks.
In cases of oversubscription, where the number of active processes exceeds the number of cores allocated to the job, both of the proposed strategies can be applied. In this context, the vector $A$ can be adjusted to reflect the expected level of oversubscription on each node. However, for the first strategy, it is necessary to ensure that all non-zero entries of $A$ are equal across nodes.
An additional important consideration appears for the initial process group that starts the application. When the execution begins using more than one node, the resulting MCW cannot returns nodes partially. Within the MaM library, there are two possible approaches to handle this situation:
  • Detect this condition during MaM initialization and recreate all processes using the Baseline method together with the parallel spawning strategy described in this work; or
  • Detect the issue and postpone any corrective action until it is actually required.
In this work, the second approach has been adopted. The reason is that the application may never need to reduce its resource usage. If a shrink is requested, the action to be taken depends on the execution history:
  • No previous expansions have been performed. The application still relies entirely on the initial MCW, which cannot be partially returned. Therefore, in case a future shrink is required, MaM performs a parallel expansion to enable TS.
  • At least one expansion has been performed. Two cases may arise depending on the requested reduction size:
    • The number of nodes to be returned is smaller than the original allocation. Only the nodes allocated in previous expansions are returned, preserving the initial MCW intact. The issue is postponed.
    • The number of nodes to be returned is greater than or equal to the original allocation. The entire initial MCW can be safely released. If additional nodes must be returned, they come from the expanded set as well. In this scenario, the challenging MCW disappears entirely.
  • The RMS requests the release of a subset of cores within a node. The excess ranks can be turned into zombie processes, enabling a partial shrink with ZS [17].
4.7 Enabling a TS shrink
To ensure a TS can be used instead of ZS, some information must be recorded in the root rank of the global communicator and in the root rank of each MCW.
The root rank of the global communicator must maintain a structure that contains for each MCW, the nodelist where they are executing. This structure enables the root to decide which approach of Subsection 4.6 should be used when shrinking and decide which groups to terminate. Additionally, if the root rank decides that its own group should be terminated: first, it selects as the new root rank the rank whose identifier is the lowest between the ones that will stay after the reconfiguration; and second it must communicate the structure to that new root rank of the global communicator.
At the MCW level, the root rank maintains a structure that records, for each rank in the communicator, whether it is active or marked as a zombie, as well as the node on which it resides. When a partial shrink affecting only a subset of cores within a node is requested, the corresponding ranks are marked as zombies and enter a sleep state. If, instead, all ranks in the MCW are marked as zombies, the group transitions to a TS operation: the zombie ranks are awakened and all processes in the communicator terminate.
If the RMS requests the release of all nodes associated with an MCW, the group is terminated directly using TS. However, if an MCW spans multiple nodes and the RMS requests the release of only a subset of those nodes, TS cannot be applied. In this case, ranks cannot be selectively marked as zombies, and the root of the global communicator must fall back to an alternative shrink strategy based on ZS.